for the function f(x) = x^2 - 2x +1
find f(1) and solve f(x) =0

Answer by edjones(8007)

f(x) = x^2 - 2x +1
.
f(0)=1
.
f(1)=1^2-2*1+1
=1-2+1
=0
.
Ed